# A Unified Dark-Matter–Driven Relativistic Bondi Route to Black-Hole Growth from Stellar to Supermassive Scales

### v1

@@ I Introduction
The discovery of luminous quasars hosting supermassive black holes (SMBHs) with masses $\sim 10^{9}M_{\odot}$ at $z\gtrsim 6$77 Banados et al. (2018); Yang et al. (2020); Larson et al. (2023a, b); Kokorev et al. (2023); Maiolino et al. (2023); Furtak et al. (2023); Harikane et al. (2023); Kocevski (2024); Kovacs et al. (2024); Bogdan et al. (2024); Juodžbalis et al. (2024); Kocevski et al. (2025); Taylor et al. (2025) presents one of the sharpest timing challenges in cosmology: assembling such masses within the first $\sim 700$ Myr strains the limits of standard formation channels based on Eddington-limited gas accretion, direct-collapse seeds, or hierarchical mergers Wise et al. (2019); Bromm et al. (2002); Oh and Haiman (2002); Begelman et al. (2006); Volonteri (2010); Inayoshi et al. (2020); Smith and Bromm (2019); Greene et al. (2020); Prole et al. (2024). Pop III remnants require prolonged, uninterrupted Eddington phases, while forming black holes from direct collapse occurs only in rare metal-poor environments. Merger-driven pathways — from classical hierarchical assembly Begelman et al. (1980); Volonteri et al. (2003) to modern cluster-driven and gravitational-wave-era hierarchical merger models Miller and Hamilton (2002); Rodriguez et al. (2019); Gerosa and Fishbach (2021) — can build stellar or intermediate-mass black holes (IMBHs), but generally lack the duty cycle and retention efficiency needed to reach $\sim 10^{9}M_{\odot}$ by $z\sim 7$ Inayoshi et al. (2020); Greene et al. (2020). These difficulties motivate additional, non-baryonic growth channels.
@@
where $M$ is the mass of the central black hole, $\Gamma=4\pi G_{N}m_{p}/\eta c\sigma_{T}$ (with the radiation efficiency $\eta\simeq 0.1$) is the baryonic Eddington coefficient, and $K$ is the coefficient for Bondi accretion. It can be determined by the equation of state (EoS) of the underlying fluid and will, in general, depend on the sound speed $a$ and the ambient mass density $\rho_{0}$, e.g., $K\sim a^{-3}\rho_{0}$ for the fluid of pressureless matter such as cold dark matter (CDM). Since dark matter (DM) prevails in the Universe and forms the main part of the typical halos, it should be responsible for feeding the central black hole in a halo via Bondi accretion. The Bondi accretion is sub-dominant when $M$ is small, and will start to dominate over the Eddington accretion at the transition mass $M_{\star}\simeq{\Gamma}/K$.
For the Bondi-aided Eddington accretion to achieve the formation of $10^{9}M_{\odot}$ SMBHs at $z\simeq 7$, it requires to have small $M_{\star}$, thus large $K$, around $z\sim 20\textrm{--}30$, during which the galaxy-scale halos with masses $\sim 10^{11}\textrm{--}10^{12}M_{\odot}$ forms, so that the central black hole can acquire most of the halo’s mass to reach $10^{9}M_{\odot}$ at $z\simeq 7$ via the dominant Bondi accretion. However, this condition will usually be destroyed by a non-negligible thermal sound speed, yielding a small Bondi radius, so that the Bondi accretion will not be complete. The above no-go results happen for CDM  [^1] See the supplementary material for more detailed discussions, unless a heavy central core, such as an IMBH, appears around $z\sim 20\textrm{--}30$, which requires specific scenarios Begelman et al. (2006); Chiu et al. (2025). This then motivates us to consider the alternative models of DM other than CDM, and the associated Bondi accretion.
@@
${\sigma\over m}\sim 0.1\text{\rm--}10~\rm{cm^{2}g^{-1}}\;.$ (2)
This simple constraint cannot fully pin down a SIDM model, leaving many possibilities open.
@@ II $\phi^{4}$-SIDM and critical Bondi accretion
where $a_{s}$ is the sound speed $a:=\sqrt{\partial p\over\partial\rho}$ at the sonic horizon, and the $\rho_{\rm DM}$ is the ambient DM mass density. We see that there is a critical limit as $a_{s}\rightarrow 1/3$, at which the Bondi accretion rate diverges. This is quite different from the Bondi accretion rate $\sim a^{-3}_{s}M^{2}\rho_{\rm DM}$ for CDM, which cannot be divergent as $a_{s}$ is non-negligible.
As our goal is to achieve the dominant Bondi accretion, we consider only the critical regime by introducing $\epsilon:=a_{s}^{2}-1/9\ll 1$. In this critical regime, we find  [^4] In fact, $\epsilon\simeq{2\over 9}{\rho_{\infty}\over\rho_{B}}$, where $\rho_{\infty}$ is the asymptotic energy density and generally differs from the mass density $\rho_{\rm DM}$ unless the asymptotic pressure $p_{\infty}$ is vanishing. In the supplementary material, we pointed out that $p_{\infty}\simeq{81\over 64}\rho_{B}\epsilon^{2}$ is one order higher in $\epsilon$ than $\rho_{\infty}\simeq{9\over 2}\rho_{B}\epsilon$. We can then treat $p_{\infty}\simeq 0$, so that $\rho_{\infty}\simeq\rho_{\rm DM}$.
@@ III Unified growth equation and transition mass
with a typical virial overdensity $\Delta_{\rm{vir}}\simeq 300$ movitavted by spherical-collapse modeling and simulations Bryan and Norman (1998); White and Rees (1978); Blumenthal et al. (1984); Schaye et al. (2015). In this environment, the Bondi radius given by (10) grows approximately linearly with black-hole mass and decreases with redshift,
@@
$m\;({\rm eV})$ $M_{\star}\;(M_{\odot})$ Growth regime
$10^{-3}$ $1.5307\times 10^{9}$ Bondi never dominates
$10^{-2}$ $4.8403\times 10^{6}$ Bondi dominates around $z\sim 11$
$10^{-1}$ $1.5307\times 10^{4}$ Early Bondi onset
11 $4.8403\times 10^{1}$ Bondi dominates almost immediately
Table 1: Transition mass $M_{\star}=\Gamma/K$ at which relativistic SIDM Bondi inflow overtakes baryonic Eddington growth. Because $K\!\propto\!m^{5/2}$, the transition scale spans nearly eight orders of magnitude across the range $m=10^{-3}$11 eV. For light SIDM, $M_{\star}\!\gtrsim\!10^{9}M_{\odot}$ so PBHs remain Eddington-limited, whereas for $m\!\gtrsim\!10^{-2}$ eV the Bondi phase switches on early, imprinting a strong microphysical signature on the black-hole mass function.
Figure 1: Mass evolution of a $10\,M_{\odot}$ seeding black hole from $z=30$ to $z=7$, including Eddington accretion and relativistic SIDM Bondi inflow. Shown are four representative SIDM masses, $m=1,~10^{-1},10^{-2}$, and $10^{-3}$ eV, evolved in a Press–Schechter median halo enhanced by an order of magnitude with a total SIDM supply fraction $f_{\rm halo}=0.9$. The heavier SIDM reaches the transition mass $M_{\star}$ (marked by crosses) earlier and experiences a rapid growth phase until saturating at the finite halo supply. In comparison, the lighter SIDM remains and experiences rapid growth at a later period.
@@
Choosing some representative values of $m$ satisfying (15), we exemplify the mass evolution of (11) in Fig. 1 for an initial mass of $10M_{\odot}$, and the Bondi radius counterpart $r_{B}$ of (13) in Fig. 2. As expected, the heavier SIDM has a sharper growth and a lower transition mass $M_{\star}$ (marked by crosses in Fig. 1), so that it is easier to yield the $10^{9}M_{\odot}$ SMBHs.
@@
Figure 2: Counterpart evolution of the Bondi radius $r_{B}(z)$ of Fig. 1, assuming the average-halo density of (12). The $y$-axis is in parsecs (logarithmic), while the $x$-axis is linear in redshift from $z=30$ (left) to $z=7$. This demonstrates how the Bondi radius expands across cosmic epochs, eventually approaching or crossing the inner-halo scale (represented by the dashed line). This enables efficient SIDM feeding in the late growth phase to form SMBHs with masses $\geq 10^{9}M_{\odot}$ if the crossing occurs at $z>7$.
@@
Due to the window of $\sigma/m$ given in (2), the SIDM develops constant-density cores through heat conduction driven by DM self-scattering, with simulations typically giving $r_{\rm core}\sim 0.02\textrm{--}0.10R_{\rm vir}$ Balberg et al. (2002); Kaplinghat et al. (2016); Tulin and Yu (2018). Here $R_{\rm{vir}}$ is the virial radius enclosing an overdensity factor $\Delta_{\rm{vir}}\sim 200\textrm{--}300$ relative to the mean matter density, providing a natural scale for the halo boundary. We overlay this core band to indicate when the Bondi radius $r_{B}(z)$ exceeds the SIDM core in Fig. 2, signaling the onset of supply-limited growth. This is the key ingredient for guaranteeing the formation of SMBHs with masses comparable to the total mass of the hosting dark halos.
@@ IV PBH seeds to mass function at $z=7$
We adopt a subdominant PBH abundance $f_{\rm{PBH}}=10^{-3}$, consistent with the limits from CMB anisotropy and spectral-distortion constraints on early energy injection ($f_{\rm{PBH}}\lesssim 10^{-3}$Ricotti et al. (2008); Ali-Haïmoud and Kamionkowski (2017); Poulin et al. (2017); Serpico et al. (2020). Independent bounds from microlensing surveys (MACHO, EROS, OGLE, Subaru/HSC) likewise rule out $M\sim 10^{-10}-10^{2}M_{\odot}$ PBHs as the dominant DM component Tisserand and others (EROS-2 Collaboration); Niikura et al. (2019); Carr and Kühnel (2020). Although QCD PBHs cannot supply the total dark matter, subdominant populations remain viable and can participate in astrophysical growth, providing a physically motivated population of early seeds.

### v2

@@ I Introduction
The discovery of luminous quasars hosting supermassive black holes (SMBHs) with masses $\sim 10^{9}M_{\odot}$ at $z\gtrsim 6-7$ Banados and others (2018); Yang and others (2020); Larson and others (2023b, a); Kokorev and others (2023); Maiolino and others (2024); Furtak and others (2023); Harikane and others (2023); Kocevski (2024); Kovacs and others (2024); Bogdan and others (2024); Juodžbalis and others (2024); Kocevski and others (2025); Taylor and others (2025) presents one of the sharpest timing challenges in cosmology: assembling such masses within the first $\sim 700$ Myr strains the limits of standard formation channels based on Eddington-limited gas accretion, direct-collapse seeds, or hierarchical mergers Wise et al. (2019); Bromm et al. (2002); Oh and Haiman (2002); Begelman et al. (2006); Volonteri (2010); Inayoshi et al. (2020); Smith and Bromm (2019); Greene et al. (2020); Prole et al. (2024). Pop III remnants require prolonged, uninterrupted Eddington phases, while forming black holes from direct collapse occurs only in rare metal-poor environments. Merger-driven pathways — from classical hierarchical assembly Begelman et al. (1980); Volonteri et al. (2003) to modern cluster-driven and gravitational-wave-era hierarchical merger models Miller and Hamilton (2002); Rodriguez et al. (2019); Gerosa and Fishbach (2021) — can build stellar or intermediate-mass black holes (IMBHs), but generally lack the duty cycle and retention efficiency needed to reach $\sim 10^{9}M_{\odot}$ by $z\sim 7$ Inayoshi et al. (2020); Greene et al. (2020). These difficulties motivate additional, non-baryonic growth channels.
@@
where $M$ is the mass of the central black hole, $\Gamma=4\pi G_{N}m_{p}/\eta c\sigma_{T}\approx 4.5\times 10^{-8}{\rm yr^{-1}}$ (assuming radiation efficiency $\eta\simeq 0.1$) is the baryonic Eddington coefficient, and $K$ is the coefficient for Bondi accretion. It can be determined by the equation of state (EoS) of the underlying fluid and will, in general, depend on the sound speed $a$ and the ambient mass density $\rho_{0}$, e.g., $K\sim a^{-3}\rho_{0}$ for the fluid of pressureless matter such as cold dark matter (CDM). Since dark matter (DM) prevails in the Universe and forms the main part of the typical halos, it should be responsible for feeding the central black hole in a halo via Bondi accretion. The Bondi accretion is sub-dominant when $M$ is small, and will start to dominate over the Eddington accretion at the transition mass $M_{\star}\simeq{\Gamma}/K$.
For the Bondi-aided Eddington accretion to achieve the formation of $10^{9}M_{\odot}$ SMBHs at $z\simeq 7$, it requires to have small $M_{\star}$, thus large $K$, around $z\sim 20-30$, during which the galaxy-scale halos with masses $\sim 10^{11}-10^{12}M_{\odot}$ forms, so that the central black hole can acquire most of the halo’s mass to reach $10^{9}M_{\odot}$ at $z\simeq 7$ via the dominant Bondi accretion. However, this condition will usually be destroyed by a non-negligible thermal sound speed, yielding a small Bondi radius, so that the Bondi accretion will not be complete. The above no-go results happen for CDM  [^1] See the supplementary material for more detailed discussions, unless a heavy central core, such as an IMBH, appears around $z\sim 20-30$, which requires specific scenarios Begelman et al. (2006); Chiu et al. (2025). This then motivates us to consider the alternative models of DM other than CDM, and the associated Bondi accretion.
@@
${\sigma\over m}\sim 0.1-1~\rm{cm^{2}g^{-1}}\;.$ (2)
This simple constraint cannot fully pin down a SIDM model, leaving many possibilities open. Here, we shall mention the pioneer works by Ostriker Ostriker (1999, 2000) by introducing the SIDM with velocity-dependent cross-section constrained by (2) to explain the formation of SMBHs via Bondi accretion and predict the Magorrian relation ($M_{\rm BH}\propto M_{\rm halo}$Magorrian and others (1998b). This approach is based on the particle picture of SIDM scattered by the Newtonian potential around a central black hole. The particle picture is intuitive and flexible in modeling the velocity dependence of the cross-section. In contrast, in this Letter, we will consider the relativistic perfect-fluid picture of a minimal SIDM that forms SMBHs via Bondi accretion. The fluid picture of SIDM captures the collective behavior of SIDM beyond the particle picture and explains the formation of SMBHs through a mechanism quite different from that in the particle picture.
@@
Despite the success of our model in explaining the SMBHs at $z\simeq 7$, our SIDM model yields a velocity-independent cross section $\sigma_{\phi^{4}}={\lambda^{2}\over 64\pi m^{2}}$ and cannot help explain the multiscale observations, known as the Goldilocks problem of dark matter halos, see e.g., Tulin and Yu (2018): galaxy-scale constraints require higher values of $\sigma/m$, but the cluster scale constraints require lower values. This implies that we need to add a new ingredient to our SIDM model to introduce velocity dependence of the cross-section to resolve the Goldilocks problem. One way is to introduce another SIDM with a velocity-dependent cross-section to solve the Goldilocks issue, but it will not contribute significantly to Bondi accretion, as is usually expected. Another way is to introduce a mediator field of mass $m_{I}$ to our $\phi^{4}$-SIDM, which can solve the Goldilocks problem as shown in, for example, Tsai et al. (2022) with $m_{I}\gtrsim{\cal O}({\rm MeV})$. On the other hand, as argued in supplementary, a heavy mediator field will not change the nature of the fluid of $\phi^{4}$-SIDM for considering the critical Bondi accretion. This can be simply seen by integrating out the heavy mediator field to restore the $\phi^{4}$-SIDM with a renormalized quartic coupling.
@@ II $\phi^{4}$-SIDM and critical Bondi accretion
where $a_{s}$ is the sound speed $a:=\sqrt{\partial p\over\partial\rho}$ at the sonic horizon, and the $\rho_{\rm DM}$ is the ambient DM mass density. This Bondi accretion rate seems divergent as $a_{s}\rightarrow 1/3$. It is quite different from the Bondi accretion rate $\sim a^{-3}_{s}M^{2}\rho_{\rm DM}$ for CDM, which cannot be divergent as $a_{s}$ is non-negligible. Moreover, as argued in supplementary and discussed right below, this seemingly divergence of the accretion rate at the critical limit $a_{s}\rightarrow 1/3$ turns out to be physically finite and environment-independent. This is the key result to be adopted for the SMBH formations.
As our goal is to achieve the dominant Bondi accretion, we consider only the critical regime by introducing $\epsilon:=a_{s}^{2}-1/9\ll 1$. In this critical regime, the consistency relations of Bondi accretion formalism yield [^4] For details, see supplementary. There, we first obtain $\epsilon\simeq{2\over 9}{\rho_{\infty}\over\rho_{B}}$, where $\rho_{\infty}$ the asymptotic energy density and generally differs from the mass density $\rho_{\rm DM}$ unless the asymptotic pressure $p_{\infty}$ is vanishing. We further pointed out that $p_{\infty}\simeq{81\over 64}\rho_{B}\epsilon^{2}$ is one order higher in $\epsilon$ than $\rho_{\infty}\simeq{9\over 2}\rho_{B}\epsilon$. We can then treat $p_{\infty}\simeq 0$, so that $\rho_{\infty}\simeq\rho_{\rm DM}$.
@@ III Unified growth equation and transition mass
with a typical virial overdensity $\Delta_{\rm{vir}}\simeq 300$ movitavted by spherical-collapse modeling and simulations Bryan and Norman (1998); White and Rees (1978); Blumenthal et al. (1984); Schaye and others (2015). In this environment, the Bondi radius given by (10) grows approximately linearly with black-hole mass and decreases with redshift,
@@
$m\;({\rm eV})$ $M_{\star}\;(M_{\odot})$ $z_{\star}$ Growth regime
$10^{-3}$ $1.5307\times 10^{9}$ none Bondi never dominates
$10^{-2}$ $4.8403\times 10^{6}$ 7.8 Bondi dominates around $z\sim 11$
$10^{-1}$ $5307\times 10^{4} 11.$31 Early Bondi onset
11 $4.8403\times 10^{1}$ 22.83 Bondi dominates almost immediately
Table 1: Transition mass $M_{\star}=\Gamma/K$ at which relativistic SIDM Bondi inflow overtakes baryonic Eddington growth, and we denote $z_{\star}$ as the corresponding cosmic redshift marking the onset of this Bondi-dominated accretion phase. Because $K\!\propto\!m^{5/2}$, the transition scale spans nearly eight orders of magnitude across the range $m=10^{-3}$11 eV. For light SIDM, $M_{\star}\!\gtrsim\!10^{9}M_{\odot}$ so PBHs remain Eddington-limited, whereas for $m\!\gtrsim\!10^{-2}$ eV the Bondi phase switches on early, imprinting a strong microphysical signature on the black-hole mass function.
Figure 1: Mass evolution of a $10\,M_{\odot}$ seeding black hole from $z=30$ to $z=7$, including Eddington accretion and relativistic SIDM Bondi inflow. Shown are four representative SIDM masses, $m=1,~10^{-1},10^{-2}$, and $10^{-3}$ eV, evolved in a Press–Schechter median halo enhanced by an order of magnitude with a total SIDM supply fraction $f_{\rm halo}=0.9$. The heavier SIDM reaches the transition mass $M_{\star}$ at the corresponding redshift $z_{\star}$, marked by crosses) earlier and experiences a rapid growth phase until saturating at the finite halo supply. These results demonstrate that the accretion process involves three stages: (i) a long early-time Eddington-dominated phase; (ii) supply-limited Bondi-dominated phase starting around $z=z_{\star}$ and ending at $z=z_{B}$ after exhausting the SIDM of the inner halo core, indicated by the sharp turns into a universal plateau; (iii) late-time Eddington-dominated phase, i.e., shown by the plateau.
@@
Choosing some representative values of $m$ satisfying (15), we exemplify the mass evolution of (11) in Fig. 1 for an initial mass of $10M_{\odot}$, and the Bondi radius counterpart $r_{B}$ of (13) in Fig. 2. As expected, the heavier SIDM has a sharper growth and a lower transition mass $M_{\star}$ (marked by crosses in Fig. 1), so that it is easier to yield the $10^{9}M_{\odot}$ SMBHs. In arriving Fig. 1 and 2, we assume that most of the SIDM resides in the inner halo core, so that the Bondi accretion is supply-limited once the inner core SIDM is accreted into the center black hole. Specifically, the Bondi-dominated phase starts at $z=z_{\star}$, where $M=M_{\star}$, and soon can engulf most of the inner core SIDM’s mass into the black hole not long after the Bondi radius $r_{B}$ exceeds $r_{\rm core}$, the inner core size of the halo. After consuming most of the SIDM in the inner core, the accretion process abruptly exits the Bondi-dominated phase at $z=z_{B}$ and re-enters the universal Eddington-dominated phase. These transitions are indicated in Fig. 1 and 2 by the sharp turns into an Eddington plateau at $z=z_{B}$. To maintain physical consistency, total growth is constrained by a mass-supply cap, which we assume to be $f_{\rm halo}=0.9$ fraction of the total halo mass, thereby restricting the final BH mass to about $90\%$ of the host halo’s total mass budget around $z=z_{B}$.
@@
Figure 2: Counterpart evolution of the Bondi radius $r_{B}(z)$ of Fig. 1, assuming the average-halo density of (12). The $y$-axis is in parsecs (logarithmic), while the $x$-axis is linear in redshift from $z=30$ (left) to $z=7$. This demonstrates how the Bondi radius expands across cosmic epochs, eventually approaching or crossing the Milky Way-like inner-halo scale ($r_{\rm{core}}\approx 0.05R_{\rm{vir}}$, represented by the gray dashed line and shaded band). This enables efficient SIDM feeding in the late growth phase to form SMBHs with masses $\geq 10^{9}M_{\odot}$ if the crossing occurs at $z>7$.
@@
Due to the window of $\sigma/m$ given in (2), the SIDM develops constant-density cores through heat conduction driven by DM self-scattering, with simulations typically giving $r_{\rm core}\sim 0.02\textrm{--}0.10R_{\rm vir}$ Balberg et al. (2002); Kaplinghat et al. (2016); Tulin and Yu (2018). This environment is modeled as a Milky Way-like dark matter core nested inside a massive host halo. We adopt a halo reference mass of $M_{\rm{ref}}\approx 5\times 10^{9}M_{\odot}$ at $z=10$, derived from the Press-Schechter (PS) formalism to represent high-density fluctuation peaks from $z=30$ to $7$. Here $R_{\rm{vir}}$ is the virial radius enclosing an overdensity factor $\Delta_{\rm{vir}}\sim 200\textrm{--}300$ relative to the mean matter density, providing a natural scale for the halo boundary. The characteristic inner core size of the host halo is modeled after a Milky Way-like profile (gray dashed line in Fig. 2), serving as a structural benchmark for the onset of supply-limited accretion. We overlay this core band to indicate when the Bondi radius $r_{B}(z)$ exceeds the SIDM core in Fig. 2, signaling the onset of supply-limited growth. This is the key ingredient for guaranteeing the formation of SMBHs with masses comparable to the total mass of the hosting dark halos.
Our results in Fig. 1 and 2 indicate that the formation of SMBHs can end around $z=z_{B}\simeq 25-10$ with their mass about $f_{\rm halo}M_{\rm halo}(z\simeq z_{B})$, earlier than the observed time around $z=7$ by the observations of the James Webb Space Telescope (JWST). The subsequent cosmic evolution of halo mass assembly through hierarchical mergers will increase halo mass but not the mass of the central SMBH. Thus, the halo mass typically scales as $(1+z)^{-\alpha}$ with $\alpha=3-4$ O’Shea et al. (2015); Griffen et al. (2016); Lacey and Cole (1994); Bond et al. (1991). This will then reduce the mass ratio $M_{\rm BH}(z_{B})/M_{\rm halo}(z)$ by a factor of $({1+z\over 1+z_{B}})^{\alpha}$. For $\alpha=3$, $z=7$ and $z_{B}=25-10$, we will obtain $M_{\rm SMBH}/M_{\rm halo}\simeq 0.3-0.03$. The order of this mass ratio is remarkably consistent with recent JWST observations of high-redshift quasars and galaxies Harikane and others (2023); Maiolino and others (2024); Bogdan and others (2024), but exceeds the local Magorrian relation ($\sim 10^{-3}$) Magorrian and others (1998a); Kormendy and Ho (2013). This indicates that the SMBHs formed in our "early blossom of SMBHs vs late-time assembly of halos" scenario are, in general, "overmassive" compared with their lower-redshift counterparts.
@@ IV PBH seeds to mass function at $z=7$
We adopt a subdominant PBH abundance $f_{\rm{PBH}}=10^{-3}$, consistent with the limits from CMB anisotropy and spectral-distortion constraints on early energy injection ($f_{\rm{PBH}}\lesssim 10^{-3}$Ricotti et al. (2008); Ali-Haïmoud and Kamionkowski (2017); Poulin et al. (2017); Serpico et al. (2020). Independent bounds from microlensing surveys (MACHO, EROS, OGLE, Subaru/HSC) likewise rule out $M\sim 10^{-10}-10^{2}M_{\odot}$ PBHs as the dominant DM component Tisserand and others (EROS-2 Collaboration) (2007); Niikura et al. (2019); Carr and Kühnel (2020). Although QCD PBHs cannot supply the total dark matter, subdominant populations remain viable and can participate in astrophysical growth, providing a physically motivated population of early seeds.